Determining missing items in return and residual income computations

Data for Uberto Company are presented in the following table of returns on investment and residual incomes:


Invested Assets

Income from Operations

Return on Investment
Minimum Return on Investment Minimum Acceptable Income from Operations
Residual Income
$870,000 $200,100 (a) 13% (b) (c)
$500,000 (d) (e) (f) $55,000 $25,000
$390,000 (g) 14% (h) $39,000 (i)
$290,000 $58,000 (j) 12% (k) (l)

Determine the missing values, identified by the letters above. For all amounts, round to the nearest whole number.

a. %
b. $
c. $
d. $
e. %
f. %
g. $
h. %
i. $
j. %
k. $
l. $
Answer #1

Determine the missing values, identified by the letters above. For all amounts, round to the nearest whole number.

a. 200100/870000 = 23 %
b. 870000*13% = $113100
c. $87000
d. $55000+25000 = 80000
e. 80000/500000 = 16 %
f. 55000/500000 = 11 %
g. $390000*14% = 54600
h. 39000/390000 = 10 %
i. $54600-39000 = 15600
j. 58000/290000 = 20 %
k. 290000*12% = 34800
l. $58000-34800 = 23200